# Operationalizing content moderation “accuracy” in the Digital Services Act

### v3

@@ Results
Based on our simulation in CivilComments, our recommendations for conducting stratified sampling in practice are: (1) Choose quantile binning. Among the two practical binning methods, quantile binning outperforms equal width binning. It performs at near oracle binning performance if the number of bins is sufficiently large and the optimal allocation is used. (2) Choose pilot allocation. Attempting to estimate a good allocation with pilot samples will greatly reduce the variance of the stratified sampling estimator. (3) Make a reasonable choice for the number of pilot samples. If the prevalence is suspected to be low, or the estimation requirement is precise, inefficiency in pilot samples is less of a concern (because the pilot samples are needed anyways). When the prevalence is high and the number of samples required may be less than the total pilot samples, a reasonable choice needs to be made about the number of pilot samples collected. The experimenter must consider how much data annotation is acceptable at worst. In Appendix C.2 we show that these recommendations are robust for lower prevalences as well.
6. Recall reporting in practice
In this section, we present a case study of recall reporting for Reddit. As the requirement currently underspecifies the technical implementation, we identify a number of ways recall could be reported. We then choose one to report and discuss the implications and need for legal clarification. We choose to focus on Reddit for two reasons: first, Reddit qualifies as an “online platform” as defined in Art. 3(i) DSA and therefore is obliged to provide transparency reports annually (Art. 15(1) DSA). [^17] https://ec.europa.eu/commission/presscorner/detail/en/ip_23_2413 Second, Reddit data is most readily available for academic study.
6.1. Addressing underspecification
@@
As discussed in §3, we establish recall as a technically and legally sensible metric, which we showed in §5 can be efficiently estimated with stratified sampling.
@@
  • Illegal content. To identify illegal content, legal definitions (e.g., punishable hate speech) would need to be operationalized as clear annotation guidelines (Zufall et al., 2022). To be feasible at the scale social media requires, such identification likely needs to be possible for a skilled person with no legal training.
  • Content that violates terms and conditions. A convenience in identifying whether speech violates community guidelines as a part of terms and conditions is that it is already operationalized by the platform. The measurement of accuracy on the guidelines is a measurement of whether the moderation system is achieving its intended purpose. This would also complement other transparency measures about the guidelines themselves (see Art. 14 DSA).
As the scope of the DSA applies to all content, and all platforms have community guidelines (Gillespie, 2018), this specification may be general enough to allow for explicit legal clarification. However, regulators should take caution when specifying recall on illegal content. Such reporting is only feasible if the respective legal bases defining content illegality are amenable to operationalization.
Test set.
@@
  • All visible content. An accuracy metric could be evaluated over all visible content on the platform, which would follow industry best practices for prevalence metrics (see §2).
  • All notified content. Art. 16 DSA requires providers to put in place mechanisms to allow users to notify (i.e., flag) content they consider to be illegal. When a notice is submitted, providers are liable if they do not act expeditiously to provide a legal examination and remove or disable access to this content if it is illegal (Art. 16(3), Art. 6(1)(b) DSA). As platforms are expected to review this content, the moderation recall of this subset may be especially important. However, anyone can flag content without concrete reason, possibly even with malicious intent (Meyer, 2014).
The specification of a test set will be heavily platform dependent. For Reddit, content may mean posts or comments, but for Youtube it may mean videos. In Snapchat, most content is not publicly visible, and it may be more appropriate to calculate recall over views. Regulators should be wary of overgeneralizing, as social media platforms have diverse functionality (Gillespie et al., 2020). Underspecifying the test set allows for custom implementation and may encourage meaningful reporting.
@@
6.2. Experimental setup
Study of interest
We restrict our study to a recent dump of Reddit (December 2022) on Pushshift.io (Baumgartner et al., 2020). We select 6 subreddits with a range of macro-norm violation prevalence, as estimated in Park et al. (2022), and that have more than 1000 comments per day [^19] Accessed January 21st, 2023, as per https://subredditstats.com/.:
Data collection
Removed comments
While publicly viewable posts from Reddit are accessible through the Pushshift data stores, records for comments removed are incomplete. Therefore, we can only derive an upper bound of content moderation recall by assuming every removal decision is a true positive.
@@ 6.2. Experimental setup
Table 1. Statistics derived from pilot annotations. Stratified sampling was applied with 8 quantile bins, 50 pilot annotations per bin, and a Roberta finetuned on personal attack data. Estimating the prevalence here is equivalent to using equal proportion allocation. The number of samples required and efficiency for estimation with random and pilot stratified sampling is provided assuming the lower bound of prevalence.
@@ Results
Refer to Table 1. As a sanity check, we see that ranking subreddits by personal attack prevalence aligns closely with their ranking by macro-norm violation as estimated in Park et al. (2022). By the prevalences, we can conclusively determine that r/politics has the biggest problem with personal attacks out of all the subreddits we examine. Moderation recall is also low for r/politics. While they have a rule against personal attacks, we speculate that moderators avoid removing too many comments to appear politically neutral. For r/legaladvice, personal attack prevalence is already low, and moderators also remove many comments that are off topic, which may inflate the true positives and consequently the recall. In terms of methodology, applying our version of stratified sampling often yields substantial sample reductions. These gains are present despite distribution shift, where the binning classifier was trained on r/changemyview but tested on other subreddits.
@@
There are two discrepancies between our experiment and the actual reporting obligation of the DSA. First, our recall is reported over a month while Reddit only needs to provide transparency reports every year. Our method can easily be adopted to a larger pool and the estimation difficulty only increases in the larger pool if the prevalence decreases. Second, the removed comments we consider could have been removed through either automated or manual means, while only recall reporting of automated tools are required by the DSA. Since we do not have records of how the comments were removed, we cannot easily provide an isolated accuracy of the automated systems, but our method is general. However, we point out that the recall presented is in line with current industry prevalence metrics, and we suggest legislators expand the scope of the requirement to an aggregate moderation accuracy.
7. Conclusion
In this work, we identified an oversight in the DSA on the reporting of “accuracy” for content moderation. To operationalize the requirement, we considered the regulatory purposes standing behind the DSA and the underlying fundamental rights which are affected. The DSA balances freedom of expression with the right to non-discrimination, and we find precision and recall to be the correct technical instantiations of these rights, where balancing is reflected in the precision-recall tradeoff. For recall, we identify the need for its efficient and unbiased statistical estimation to respect the right to conduct business. Stratified sampling is a natural solution and we provide multiple technical contributions in its application to estimate recall. Finally, beyond the quantitative results, we provide qualitative observations of where the law underspecifies the experimental design in accuracy reporting.
8. Ethical considerations
@@
Content moderation can violate the right to the freedom of expression. Automated content moderation further raises ethical concerns about problematic training data or application of such systems perpetuating societal power dynamics (Gillespie, 2020). We do not approve the use of these systems for censorship and have especially highlighted that automated content moderation needs to respect freedom of speech. We promote the use of stratified sampling: it involves humans and is unbiased with respect to a human defined ground truth. Finally, our work focuses on the evaluation of content moderation rather than directly making removal decisions, and we hope it can help further the study of fair interventions to reduce social media harms.
@@
Regulation can be co-opted to serve minority interests. Facebook, for instance, has openly requested for additional social media regulation, which is speculated to serve their interest in preventing competition (Macias, 2020). Our position on regulation is not neutral either, as this work proposes new forms of regulation. However, the entirety of this work is dedicated to content moderation recall’s legal and technical sensibility. A large focus of our work is on technical feasibility, which is especially important in considering the compliance of new companies and preventing regulatory capture.
@@ Appendix A Additional statistical background
A.1. Formulating prevalence estimation
@@
Table 2. Number of samples needed for the random sampling estimator to estimate different prevalences, to different levels of precision. A normal approximation is assumed. The smaller the prevalence the more difficult it is to estimate precisely. Highlighted in bold is the prevalence of positives in the CivilComments dataset.
A.2. Stratified sampling notation
@@
(5) $\hat{p}_{st}$ $=\sum_{h=1}^{L}\frac{N_{h}}{N}\cdot\hat{p}_{h}$
(6) $\text{where}~{}~{}\hat{p}_{h}$ $=\frac{1}{n_{h}}\sum_{n_{h}}^{i=1}s(x^{h}_{i})\,\forall\,h\in\{1,\dotsc,L\};$
@@
(7) $\widehat{\text{Var}}(\hat{p}_{st})=\sum_{h=1}^{L}\left(\frac{N_{h}}{N}\right)^{2}\left(1-\frac{n_{h}}{N_{h}}\right)^{2}\cdot\frac{\hat{p}_{h}(1-\hat{p}_{h})}{n_{h}}.$
@@
A.3. Optimal allocation for stratified sampling
@@
(8) $n^{opt}_{h}=n\cdot\frac{N_{h}\sigma_{h}}{\sum_{k}N_{k}\sigma_{k}}$
@@
A.4. Power calculations for random sampling
@@
(9) $\textsc{SE}_{req}^{2}=\frac{\hat{p}(1-\hat{p})}{n}\qquad\therefore n=\frac{\hat{p}(1-\hat{p})}{\textsc{SE}_{req}^{2}}$
@@
A.5. Power calculations for stratified sampling
@@
(10) $\textsc{SE}_{req}^{2}$ $=\sum_{h=1}^{L}\left(\frac{N_{h}}{N}\right)^{2}\cdot\frac{\hat{p}_{h}(1-\hat{p}_{h})}{(1/L)n}$
(11) $\therefore n$ $=\left(\frac{L}{\textsc{SE}_{req}^{2}}\right)\sum_{h=1}^{L}\left(\frac{N_{h}}{N}\right)^{2}\cdot\hat{p}_{h}(1-\hat{p}_{h})$
@@
(12) $\textsc{SE}_{req}^{2}$ $=\sum_{h=1}^{L}\left(\frac{N_{h}}{N}\right)^{2}\cdot\frac{\hat{p}_{h}(1-\hat{p}_{h})}{n^{opt}_{h}}$
(13) $\therefore n$ $=\left(\frac{1}{\textsc{SE}_{req}^{2}}\right)\sum_{h=1}^{L}\left(\frac{N_{h}}{N}\right)^{2}\cdot\frac{\hat{p}_{h}(1-\hat{p}_{h})}{c^{opt}_{h}}$
@@ Appendix B Estimating precision
The qualitative difficulty is in implementing the scoring function $s(x)$. Since the ground truth will be applied to content detected by an automated system, many false positives may be borderline. In the case where the ground truth is illegal content, a trained legal expert may be needed, and some borderline content may even need to wait for judicial decisions. In the case where the ground truth are community guidelines, the decision boundaries of these guidelines are continually refined, and the moderators judging content violation need substantial training (Newton, 2019).
@@
Roberta (balanced) oracle:8 optimal 578 2312 9247
@@
Table 3. Number of samples needed for the stratified sampling estimator to estimate prevalence of toxicity (4.1%) in the unfiltered CivilComments within 5, 10, and 20%. The intervals denote the confidence intervals on the recall. Oracle:8 means 8 strata were created and pilot:50 means 50 pilot samples are taken within each strata for variance estimation. Oracle binning and optimal allocation are oracle techniques which rely on labelled data that is not available beforehand in practice. We chose 8 bins and 50 pilot samples because these settings were effective in Figure 1.
@@ Appendix C Additional stratified sampling experiments
C.1. Stratification with different classifiers

### v5

@@ Results
Pilot sampling under different prevalences
We examine the robustness of pilot sampling to different class imbalances in Figure 2. For each example, civilcomments has a real valued toxicity label corresponding to the proportion of annotators that labelled it as toxic. When converting the labels to binary, a threshold of 0.5 yields a prevalence of 5.9%, which is the setting studied in Figure 1. In Figure 2 we present results for different prevalences by varying the threshold among $\{0.48,0.83,0.95\}$. Stratified sampling provides significant efficiency over random sampling. The number of pilot samples matters for both very small and very large prevalences. In the case of high prevalence, pilot samples outnumber the optimal allocation so there will be wasted samples. In the case of low prevalence, additional pilot samples help estimate the standard deviation of the strata correctly for optimal allocation. Finally, increasing the number of strata does not help with lower prevalences, as fewer samples are available for each bin to estimate the standard deviation with.
@@
Based on our simulation in CivilComments, our recommendations for conducting stratified sampling in practice are: (1) Choose quantile binning. Among the two practical binning methods, quantile binning outperforms equal width binning. It performs at near oracle binning performance if the number of bins is sufficiently large and the optimal allocation is used. (2) Choose pilot allocation. Attempting to estimate a good allocation with pilot samples will greatly reduce the variance of the stratified sampling estimator. (3) Make a reasonable choice for the number of pilot samples. If the prevalence is suspected to be low, or the estimation requirement is precise, inefficiency in pilot samples is less of a concern (because the pilot samples are needed anyways). When the prevalence is high and the number of samples required may be less than the total pilot samples, a reasonable choice needs to be made about the number of pilot samples collected. The experimenter must consider how much data annotation is acceptable at worst.
6 Recall reporting in practice
In this section, we present a case study of recall reporting for Reddit. As the requirement currently underspecifies the technical implementation, we identify a number of ways recall could be reported. We then choose one to report and discuss the implications and areas for legal clarification. We choose to focus on Reddit for two reasons: first, Reddit qualifies as an “online platform” as defined in Art. 3(i) DSA and therefore is obliged to provide transparency reports annually (Art. 15(1) DSA). [^20] https://ec.europa.eu/commission/presscorner/detail/en/ip˙23˙2413 Second, Reddit data is most readily available for academic study.
Figure 2: Number of samples required to estimate the prevalence of toxicity within 20%, for different prevalences. All pilot results are averaged over 100 trials. The numbers associated with the pilot is the total number of pilot samples across all bins. By choosing appropriate powers of two, each bin always has an integer number of pilot samples. Dotted lined represents random sampling. For small prevalences, more pilot samples are better, because more samples estimate the standard deviations within each bin better. In contrast, more bins fares worse because each bin has less pilot samples.
6.1 Addressing underspecification
@@
As discussed in §3, we establish recall as a legally sensible metric, which we showed in §5 can be efficiently estimated with stratified sampling.
@@
  • Illegal content. To identify illegal content, legal definitions (e.g., punishable hate speech) would need to be operationalized as clear annotation guidelines (Zufall et al. 2022). To be feasible at the scale social media requires, such identification likely needs to be possible for a skilled person with no legal training.
  • Content that violates terms and conditions. A convenience in identifying whether speech violates community guidelines as a part of terms and conditions is that it is already operationalized by the platform. The measurement of accuracy on the guidelines is a measurement of whether the moderation system is achieving its intended purpose. This would also complement transparency measures about the guidelines themselves (see Art. 14 DSA).
As the scope of the DSA applies to all content, and all platforms have community guidelines (Gillespie 2018), this specification may be general enough to allow for explicit legal clarification. However, regulators should take caution when specifying recall on illegal content. Such reporting is only feasible if the respective legal bases defining content illegality are amenable to operationalization.
Test set
@@
  • All visible content. An accuracy metric could be evaluated over all visible content on the platform, which would also follow industry best practices on how prevalence metrics are reported (see §2).
  • All notified content. Art. 16 DSA requires providers to put in place mechanisms to allow users to notify (i.e., flag) content they consider to be illegal. When a notice is submitted, providers are liable if they do not act expeditiously to provide a legal examination and remove or disable access to this content if it is illegal (Art. 16(3), Art. 6(1)(b) DSA). As platforms are expected to review this content, the moderation recall of this subset may be especially important. However, anyone can flag content without concrete reason, possibly even with malicious intent (Meyer 2014).
The specification of a test set will be heavily platform dependent. For Reddit, content may mean posts or comments, but for Youtube it may mean videos. In Snapchat, most content is not publicly visible, and it may be more appropriate to calculate recall over views. Regulators should be wary of overgeneralizing, as social media platforms have diverse functionality (Gillespie et al. 2020). Underspecifying the test set allows for custom implementation and may encourage meaningful reporting.
@@
6.2 Experimental setup
@@
Table 1: Statistics derived from pilot annotations. Stratified sampling was applied with 8 quantile bins, 50 pilot annotations per bin, and a Roberta finetuned on personal attack data. Estimating the prevalence here is equivalent to using equal proportion allocation. The number of samples required and efficiency for estimation with random and pilot stratified sampling is provided assuming the lower bound of prevalence.
Study of interest
We restrict our study to a recent dump of Reddit (December 2022) on Pushshift.io (Baumgartner et al. 2020). We select 6 subreddits with a range of macro-norm violation prevalence, as estimated in Park, Seering, and Bernstein (2022), and that have more than 1000 comments per day [^22] Accessed January 21st, 2023, as per https://subredditstats.com/.:
Data collection
Removed comments
While publicly viewable posts from Reddit are accessible through the Pushshift data stores, records for removed comments are incomplete. Therefore, we can only derive an upper bound of content moderation recall by assuming every removal decision is a true positive.
@@ Results
Refer to Table 1. As a sanity check, we see that ranking subreddits by personal attack prevalence aligns closely with their ranking by macro-norm violation as estimated in Park, Seering, and Bernstein (2022). By the prevalences, we can conclusively determine that r/politics has the biggest problem with personal attacks out of all the subreddits we examine. By moderation recall, r/politics is also low. While they have a rule against personal attacks, we speculate that moderators avoid removing too many comments to appear politically neutral. For r/legaladvice, personal attack prevalence is already low, and moderators also remove many comments that are off topic, which may inflate the true positives and consequently the recall. In terms of methodology, applying our version of stratified sampling often yields substantial sample reductions. These gains are present despite distribution shift, where the binning classifier was trained on r/changemyview but tested on other subreddits.
@@
There are two discrepancies between our experiment and the actual reporting obligation of the DSA. First, our recall is reported over a month while Reddit only needs to provide transparency reports every year. Our method can easily be adopted to a larger pool and the estimation difficulty only increases in the larger pool if the prevalence decreases. Second, the removed comments we consider could have been removed through either automated or manual means, while only recall reporting of automated tools are required by the DSA. Since we do not have records of how the comments were removed, we cannot easily provide an isolated accuracy of the automated systems, but our method is general. However, we point out that the recall presented is in line with current industry prevalence metrics, and we additionally suggest legislators could expand the scope of the requirement to an aggregate moderation accuracy.
7 Conclusion
In this work, we identified “accuracy” in the DSA as an open-textured term, which requires further interpretation. To operationalize the requirement, we considered the regulatory purposes standing behind the DSA and the underlying fundamental rights which are affected. The DSA balances freedom of expression with the right to non-discrimination, and we find precision and recall to be an appropriate instantiation of these rights, where balancing is reflected in the precision-recall tradeoff. We derive an efficient and robust estimator, which we use to compute a moderation recall for Reddit. By providing a legally and technically sound interpretation of “accuracy”, we highlight the implications and areas for further legal clarification. In light of the insights here, we hope regulators can provide further legal or regulatory clarification to prevent enforcement deficits and better fulfill the regulatory purpose of the DSA.
8 Ethical considerations
@@
Content moderation can violate the right to the freedom of expression. Automated content moderation further raises ethical concerns about problematic training data or application of such systems perpetuating societal power dynamics (Gillespie 2020). We do not approve the use of these systems for censorship and have especially highlighted that automated content moderation needs to respect freedom of speech. We promote the use of stratified sampling: it involves humans and is unbiased with respect to a human defined ground truth. Finally, our work focuses on the evaluation of content moderation rather than directly making removal decisions, and we hope it can help further the study of fair interventions to reduce social media harms.
@@
Regulation can be co-opted to serve minority interests. Facebook, for instance, has openly requested for additional social media regulation, which is speculated to serve their interest in preventing competition (Macias 2020). Our position on regulation is not neutral either, as this work is rooted in existing regulation. However, the entirety of this work is dedicated to content moderation recall’s legal and technical sensibility. Our proposals introduce a meaningful transparency requirement which introduces strictly more reporting burden. A large portion of this work is dedicated to technical feasibility, which is especially important in considering the compliance of new companies and preventing regulatory capture.
@@
Acknowledgements
Discussion with Sagnik Ghosh of the Integrity Institute inspired the scope of our problem. Swabha Swayamdipta and the USC NLP group provided feedback on early versions of this work. Over several conference submissions, many anonymous reviewers helped improve the quality and clarity of this work. We thank all who have made our work possible. This work was funded by grants from Open Philanthropy, Cisco Research, and Google Research.
@@ Appendix A Additional statistical background
A.1 Formulating prevalence estimation
@@
Table 2: Number of samples needed for the random sampling estimator to estimate different prevalences, to different levels of precision. A normal approximation is assumed. The smaller the prevalence the more difficult it is to estimate precisely. Highlighted in bold is the prevalence of positives in the CivilComments dataset.
A.2 Stratified sampling notation
@@
$\hat{p}_{st}$ $=\sum_{h=1}^{L}\frac{N_{h}}{N}\cdot\hat{p}_{h}$ (5)
$\text{where}~{}~{}\hat{p}_{h}$ $=\frac{1}{n_{h}}\sum_{n_{h}}^{i=1}s(x^{h}_{i})\,\forall\,h\in\{1,\dotsc,L\};$ (6)
@@
$\widehat{\text{Var}}(\hat{p}_{st})=\sum_{h=1}^{L}\left(\frac{N_{h}}{N}\right)^{2}\left(1-\frac{n_{h}}{N_{h}}\right)^{2}\cdot\frac{\hat{p}_{h}(1-\hat{p}_{h})}{n_{h}}.$ (7)
@@
A.3 Optimal allocation for stratified sampling
@@
$n^{opt}_{h}=n\cdot\frac{N_{h}\sigma_{h}}{\sum_{k}N_{k}\sigma_{k}}$ (8)
@@
A.4 Power calculations for random sampling
@@
$\textsc{SE}_{req}^{2}=\frac{\hat{p}(1-\hat{p})}{n}\qquad\therefore n=\frac{\hat{p}(1-\hat{p})}{\textsc{SE}_{req}^{2}}$ (9)
@@
A.5 Power calculations for stratified sampling
@@
$\textsc{SE}_{req}^{2}$ $=\sum_{h=1}^{L}\left(\frac{N_{h}}{N}\right)^{2}\cdot\frac{\hat{p}_{h}(1-\hat{p}_{h})}{(1/L)n}$ (10)
$\therefore n$ $=\left(\frac{L}{\textsc{SE}_{req}^{2}}\right)\sum_{h=1}^{L}\left(\frac{N_{h}}{N}\right)^{2}\cdot\hat{p}_{h}(1-\hat{p}_{h})$ (11)
@@
$\textsc{SE}_{req}^{2}$ $=\sum_{h=1}^{L}\left(\frac{N_{h}}{N}\right)^{2}\cdot\frac{\hat{p}_{h}(1-\hat{p}_{h})}{n^{opt}_{h}}$ (12)
$\therefore n$ $=\left(\frac{1}{\textsc{SE}_{req}^{2}}\right)\sum_{h=1}^{L}\left(\frac{N_{h}}{N}\right)^{2}\cdot\frac{\hat{p}_{h}(1-\hat{p}_{h})}{c^{opt}_{h}}$ (13)
@@ Appendix B Estimating precision
The qualitative difficulty is in implementing the scoring function $s(x)$. Since the ground truth will be applied to content detected by an automated system, many false positives may be borderline. In the case where the ground truth is illegal content, a trained legal expert may be needed, and some borderline content may even need to wait for judicial decisions. In the case where the ground truth are community guidelines, the decision boundaries of these guidelines are continually refined, and the moderators judging content violation need substantial training (Newton 2019).
@@
Roberta (balanced) oracle:8 optimal 353 1410 5639
@@
Table 3: Number of samples needed for the stratified sampling estimator to estimate prevalence of toxicity (4.1%) in the unfiltered CivilComments within 5, 10, and 20%. The intervals denote the confidence intervals on the recall. Oracle:8 means 8 strata were created and pilot:50 means 50 pilot samples are taken within each strata for variance estimation. Oracle binning and optimal allocation are oracle techniques which rely on labelled data that is not available beforehand in practice. We chose 8 bins and 50 pilot samples because these settings were effective in Figure 1.
@@ Appendix C Additional stratified sampling experiments
C.1 Stratification with different classifiers